$A$ bag $A$ contains $2$ white and $3$ red balls and bag $B$ contains $4$ white and $5$ red balls. One ball is drawn at random from a randomly chosen bag and is found to be red. The probability that it was drawn from bag $B$ is:

  • A
    $\frac{5}{14}$
  • B
    $\frac{5}{16}$
  • C
    $\frac{5}{18}$
  • D
    $\frac{25}{52}$

Explore More

Similar Questions

The chance of India winning the toss is $3/4$. If it wins the toss,then its chance of victory is $4/5$,otherwise it is only $1/2$. Then the chance of India's victory is

An urn contains $25$ balls numbered $1$ to $25$. Suppose an odd number is considered a 'success'. If $2$ balls are drawn from the urn with replacement,find the probability of getting at least $2$ successes.

The probability that a man will be alive $25$ years hence is $0.3$ and the probability that his wife will be alive $25$ years hence is $0.4$. Find the probability that $25$ years hence,both will be alive.

If $P(A_1 \cup A_2) = 1 - P(A_1^c) P(A_2^c)$,where $c$ stands for complement,then the events $A_1$ and $A_2$ are

If four persons are chosen at random from a group of $3$ men,$2$ women,and $4$ children,then the probability that exactly two of them are children is:

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo